A Complete Bibliography of the Journal of C Language Translation

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

19 September 2010 Version 1.28

Title word cross-reference

- [Jae89k].

64-bit [Mac90b].

Abstract [Fie92]. Addendum [Pla93b]. Adding [Mac89a]. Addresses [Pla89c]. Aliasing [Mac89b]. Alternatives [Bro92c]. ANSI [Bro89a]. ANSI/ISO [Bro89a, Bro89b]. Architectures [Mac91c]. Arithmetic [Mac89a]. Arrays [Mac89d]. Asynchrony [Cha94b]. Author [Jae89a]. Avoiding [McK93a].

Barely [Pla90d]. Based [Jer91]. bit [Mac90b]. BLAS [DCHP92].

Books [Ano89]. Building [Par90]. Byte [Pla94d].

C [Jae89j, Pla89b]. C-Based [Jer91]. C/C [Koh91a]. Calling [Cha94b]. Case [Par90]. Cb [McK91a]. Changes [Pla89a]. Character [Goo90a]. Characters [Pla91d]. Checking [Spi93]. Class [BPRW94]. Code [Hof93]. Collector [Boe92]. Communication [Mac91b]. Compare [Din94]. Compatibility [Koh91a]. Compatible [Jaf90]. Compilation [Boe92]. Compiler [McK91b]. Compilers [MAA90]. Complex [Mac89a]. Compound [Pro93]. Computers [GL94]. Computing [Tyd92]. Concurrent [Lal90]. Conformance [Mar89].

Constant [McK93a]. Contents [Ano94]. Control [Pla90c]. Conventions [Cha94b]. Corrigendum [Ano93a]. CPP [Par90]. Cray [Mac90b]. Cumulative [Ano92].

Data [Mac91c]. Debuggers [Cha94b]. December [Jer92c]. Declarations [Spi93]. Design [Goo89]. Designated [Pro93]. Diagnosing [Pla89a]. Digital [Lea90]. Distribution [Mac91c]. DPCE [Sta94a]. Dropped [Din94].

Electronic [Jae89b]. Emitting [Lon91]. European [Mar89]. Evaluation [McK93a]. exception [Cha94a]. Exceptions [McK93a]. Execution [Mac92]. Expression [McK93a]. Expressions [Bro89d]. Extended [Bro90d]. Extending [Tyd92]. Extension [Kur91a]. Extensions [Jae89i]. Extraction [Jae89o].

Facility [Gim92]. Finalizers [Lei91]. First [Sak92]. flat [McK91a]. Floating [Mac90c]. Floating-Point [Mac90c]. Formed [Jae89i]. FORTRAN [Goo89]. Forum [Bro89c].

Garbage [Boe92]. Garbage-Collector-Safe [Boe92]. Generalizing [Pla90a]. Graphical [Lon95]. Group [Jae89i, Jae89j]. Guidelines [Jae89a].

Handling [Ste90]. Heap [Pla94a]. high [Hof93]. high-level [Hof93].

IEEE [Mac90c]. II [Jae89k]. III [Lei89a]. Imaginary [TC94]. Implementation [Hol89]. Implementing [Pla91c]. Incremental [MA91]. Independent [MAA90]. Index [Ano92]. Initializers [Mac90d]. Interfacing [All90]. Interlanguage [Mac91b]. Intermediate [Egg91].

Internationalization [Pla92b]. Internationalizing [Pla89b]. Interpretations [Bro90a]. Interpreting [Pla93a]. Iostreams [Pla93d]. ISO [Bro89a]. Issues [Mac89b]. Iterators [Pen91]. IV [Lei90].

J. [Jae89k]. Joint [Sak93a].

Languages [Hof93]. Leichter [Jae89k]. Length [Mac89d]. level [McK91a]. Library [Pla90c]. Linda [Lei89a, Lei89b]. Link [Spi93]. Literals [Pro93]. Locales [Pla91c]. Low [McK91a]. Low-level [McK91a].

Management [BPRW94].

Managing [Pla94a]. Manipulation [Goo90a]. Massively [Mac91c]. Math [Pla94b]. Meeting [Bro89a]. Meetings [Sak92]. Meets [Lei89b]. Microsoft [All90]. Miscellanea [Jae89d]. Model [Lon95]. Multi [Lal90]. Multi-Processing [Lal90]. Multibyte [Bro90d]. Multipass [McK91b].

Name [Mel89]. NCEG [Mac91d].

Normative [Pla93b]. Number [Jae89b]. Numerical [Mac89c, Jae89j]. Numerics [BPRW93].

Object [BPRW93].

Object-Oriented [BPRW93].

Objects [BC94]. Optimizers [Cha94b]. Oriented [BPRW93]. Overcoats [Pla92c]. Overflow [McK93b].

Parallel [Lei89b]. Parallelism [Pla90b]. Parasol [Jer91]. Parser [MAA90]. Parser-Independent [MAA90]. Parsing [Mac90d]. Part [Sak93b, Lei89b]. Pascal [Pla89d]. Placement [BC94]. Point [Mac90c]. Pointers [Pla89c]. Points [Pla90b]. Pollution [Mel89].

Pragmania [Jae89l]. Precisions [Pla94b]. Preprocessing [MA92a]. Primitives [Pla91b]. Processing [Hol89]. Programming [Mac89c]. Progress [Mac91d]. Project [Len89]. Proposal [Boe92]. Proposed [Lei91]. Prototype [Jae89o]. Publications [Ano89].

Qualifiers [Bro89c]. Quiet [Pla89a].

Regarding [TC94]. Report [Bro89a]. Resolving [McK91b]. Rest [Sak93b]. Restricted [Mac90e]. Retargetable [GL94]. Reusable [MA91]. Revising [Pla94c]. Revisited [Mac90a]. Routines [DCHP92].

Safe [Boe92]. Scanning [MA91]. Search [Bro92c]. Sequence [Pla90b]. Service [Mar89]. Sets [Pla91a]. Shared [BC94]. Signal [Lea90]. Size [Rit90]. Solution [Mel89]. Source [Lon91]. Space [Mel89]. Specification [DCHP92]. Standard [Jaf90, Pla90c]. Standardization [Len89]. Standardizing [Sak92]. Standards [Bro89c, Gwy93c]. Status [Jae89j]. Strong [Gim92]. Study [Par90]. Subset [McK91a]. Suite [Jae89p]. Superscalar [GL94]. Superset [GL94]. Support [Bro90d]. Survey [Jae89b]. Syntax [Fie92]. System [Lon95]. Systems [Jer91].

Table [Ano94]. Target [Hof93]. Templates [Pla95]. Temporaries [BPRW94]. Testing [Mar89]. Time [Spi93]. Tool [Jae89o]. Translating [Pla89d]. Translator [Goo89]. Transputer [Kur91b]. Trigraphs [Bro92c]. Turtles [Pla92c]. Type [Bro89c]. Typedefs [McK91b]. Types [TC94]. Typing [Gim92].

Understanding [Bro89d].

Universal [PM92]. Update [Bro90c].

Using [Hof93].

Validation [Jae89p]. Variable [Mac89d]. Variable-Size [Rit90]. Vector [GL94]. via [Hom93]. VLAs [Mac90a]. Vtbl [BC94].

Wide [Pla93d]. Works [Pla90d].

World [Mac90b].

Year [Sak93a].

References

Allison:JCLT-1-4-300

[All90] Bob Allison. Interfacing Microsoft C and FORTRAN. The Journal of C Language Translation, 1(4):300– 304, March 1990. ISSN 1042-5721.

Anonymous:JCLT-1-2-68

[Ano89] Anonymous. Books and publications. The Journal of C Language Translation, 1(2):68–72, September 1989. ISSN 1042-5721.

Anonymous:JCLT-3-4-341

[Ano92] Anonymous. Cumulative index. The Journal of C Language Translation, 3(4):341–348, March 1992. ISSN 1042-5721.

Anonymous:JCLT-5-1-60

[Ano93a] Anonymous. Corrigendum. The Journal of C Language Translation, 5(1):60, September 1993. ISSN 1042-5721.

Anonymous:JCLT-4-3-223

[Ano93b] Anonymous. Floating-point C extensions. The Journal of C Language Translation, 4(3):223–237, March 1993. ISSN 1042-5721.

Anonymous:JCLT-5-4-256

[Ano94] Anonymous. Cumulative table of contents. The Journal of C Language Translation, 5(4):256–260, June 1994. ISSN 1042-5721.

Banerji:JCLT-6-1-44

[BC94] Arindam Banerji and David L.

Cohn. Shared C++ objects and Vtbl placement. The Journal of C Language Translation, 6(1):44–60, September 1994. ISSN 1042-5721.

Boehm:JCLT-4-2-126

[Boe92] Hans-J. Boehm. A proposal for garbage-collector-safe C compilation. The Journal of C Language Translation, 4(2):126–141, December 1992. ISSN 1042-5721.

Budge:JCLT-5-1-32

[BPRW93] Kent G. Budge, James S. Peery, Allen C. Robinson, and Michael K. Wong. C++ and object-oriented numerics. The Journal of C Language Translation, 5(1):32–56, September 1993. ISSN 1042-5721.

Budge:JCLT-6-2-100

[BPRW94] Kent G. Budge, James S. Peery, Allen C. Robinson, and Michael K. Wong. Management of class temporaries in C++ translation systems. The Journal of C Language Translation, 6(2):100–111, December 1994. ISSN 1042-5721.

Brodie:JCLT-1-1-17

[Bro89a] Jim Brodie. ANSI/ISO meeting report. The Journal of C Language Translation, 1(1):17–19, June 1989. ISSN 1042-5721.

Brodie:JCLT-1-2-122

[Bro89b] Jim Brodie. ANSI/ISO status report. The Journal of C Language Translation, 1(2):122–126, September 1989. ISSN 1042-5721.

Brodie:JCLT-1-1-48

[Bro89c] Jim Brodie. Standards forum: Type qualifiers. The Journal of C Language Translation, 1(1):48–54, June 1989. ISSN 1042-5721.

Brodie:JCLT-1-3-205

[Bro89d] Jim Brodie. Understanding expressions in C. The Journal of C Language Translation, 1(3):205– 214, December 1989. ISSN 1042-5721.

Brodie:JCLT-2-1-48

[Bro90a] Jim Brodie. ANSI C interpretations report. The Journal of C Language Translation, 2(1):48–56, June 1990. ISSN 1042-5721.

Brodie:JCLT-2-3-207

[Bro90b] Jim Brodie. ANSI C interpretations report. The Journal of C Language Translation, 2(3):207– 215, December 1990. ISSN 1042-5721.

Brodie:JCLT-1-4-255

[Bro90c] Jim Brodie. C standards update. The Journal of C Language Translation, 1(4):255–263, March 1990. ISSN 1042-5721.

Brodie:JCLT-2-2-133

[Bro90d] Jim Brodie. Extended multibyte support. The Journal of C Language Translation, 2(2):133–140, September 1990. ISSN 1042-5721.

Brodie:JCLT-2-4-287

[Bro91a] Jim Brodie. ANSI C interpretations report. The Journal of C Language Translation, 2(4):287– 295, March 1991. ISSN 1042-5721.

Brodie:JCLT-3-1-42

[Bro91b] Jim Brodie. ANSI C interpretations report. The Journal of C Language Translation, 3(1):42–50, June 1991. ISSN 1042-5721.

Brodie:JCLT-3-3-206

[Bro91c] Jim Brodie. ANSI C interpretations report. The Journal of C Language Translation, 3(3):206– 213, December 1991. ISSN 1042-5721.

Brodie:JCLT-3-2-156

[Bro91d] Jim Brodie. C standards update.

The Journal of C Language Translation, 3(2):156–161, September 1991. ISSN 1042-5721.

Brodie:JCLT-4-1-9

[Bro92a] Jim Brodie. ANSI C interpretations report. The Journal of C Language Translation, 4(1):9–16, September 1992. ISSN 1042-5721.

Brodie:JCLT-4-2-80

[Bro92b] Jim Brodie. ANSI C interpretations report. The Journal of C Language Translation, 4(2):80–86, December 1992. ISSN 1042-5721.

Brodie:JCLT-3-4-310

[Bro92c] Jim Brodie. Trigraphs: The search for alternatives. The Journal of C Language Translation, 3(4):310– 318, March 1992. ISSN 1042-5721.

Chase:JCLT-5-4-229

[Cha94a] David Chase. Implementation of exception handling, Part I. The Journal of C Language Translation, 5(4):229–240, June 1994. ISSN 1042-5721.

Chase:JCLT-6-1-20

[Cha94b] David Chase. Implementation of exception handling, Part II: Calling conventions, asynchrony, optimizers, and debuggers. The Journal of C Language Translation, 6 (1):20–32, September 1994. ISSN 1042-5721.

Datardina:JCLT-3-4-295

[DCHP92] S. P. Datardina, J. J. Du Croz, S. J. Hammarling, and M. W. Pont. A proposed specification of BLAS routines in C. The Journal of C Language Translation, 3(4): 295–309, March 1992. ISSN 1042-5721.

Dinesh:JCLT-6-2-112

[Din94] T. B. Dinesh. Why compare dropped C++. The Journal of C Language Translation, 6(2):112– 136, December 1994. ISSN 1042-5721.

Eggert:JCLT-3-1-51

[Egg91] Paul Eggert. C as an intermediate language: A case study. The Journal of C Language Translation, 3 (1):51–61, June 1991. ISSN 1042-5721.

Fields:JCLT-4-1-42

[Fie92] Bob Fields. An abstract syntax for C. The Journal of C Language Translation, 4(1):42–52, September 1992. ISSN 1042-5721.

Gimpel:JCLT-3-4-259

[Gim92] Jim Gimpel. A strong typing facility for C. The Journal of C Language Translation, 3(4):259–274, March 1992. ISSN 1042-5721.

Gaissaryan:JCLT-5-3-183

[GL94] Sergey Gaissaryan and Alexey Lastovetsky. An ANSI C superset for vector and superscalar computers and its retargetable compiler. The Journal of C Language Translation, 5(3):183–198, March 1994. ISSN 1042-5721.

Goodman:JCLT-1-3-166

[Goo89] Goodman. Design of a FORTRAN to C translator, Part I. The Journal of C Language Translation, 1 (3):166–179, December 1989. ISSN 1042-5721.

Goodman:JCLT-2-2-141

[Goo90a] Fred Goodman. FORTRAN to C: Character manipulation. The Journal of C Language Translation, 2(2):141–151, September 1990. ISSN 1042-5721.

Goodman:JCLT-2-1-29

[Goo90b] Fred Goodman. FORTRAN to C:

Numerical issues. The Journal of C Language Translation, 2(1):29– 44, June 1990. ISSN 1042-5721.

Goodman:JCLT-1-4-264

[Goo90c] Fred Goodman. A FORTRAN to C translator, Part II. The Journal of C Language Translation, 1(4): 264–274, March 1990. ISSN 1042-5721.

Gwyn:JCLT-4-3-185

[Gwy93a] Douglas A. Gwyn. C standards update. The Journal of C Language Translation, 4(3):185–190, March 1993. ISSN 1042-5721.

Gwyn:JCLT-4-4-277

[Gwy93b] Douglas A. Gwyn. C standards update. The Journal of C Language Translation, 4(4):277–283, June 1993. ISSN 1042-5721.

Gwyn:JCLT-5-1-18

[Gwy93c] Douglas A. Gwyn. C Standards update. The Journal of C Language Translation, 5(1):18–21, September 1993. ISSN 1042-5721.

Gwyn:JCLT-5-2-91

[Gwy93d] Douglas A. Gwyn. C standards update. The Journal of C Language Translation, 5(2):91–94, December 1993. ISSN 1042-5721.

Gwyn:JCLT-5-3-163

[Gwy94a] Douglas A. Gwyn. C standards update. The Journal of C Language Translation, 5(3):163–174, March 1994. ISSN 1042-5721.

Gwyn:JCLT-5-4-241

[Gwy94b] Douglas A. Gwyn. C standards update. The Journal of C Language Translation, 5(4):241–245, June 1994. ISSN 1042-5721.

Gwyn:JCLT-6-1-33

[Gwy94c] Douglas A. Gwyn. C standards update. The Journal of C Language Translation, 6(1):33–34, September 1994. ISSN 1042-5721.

Gwyn:JCLT-6-2-86

[Gwy94d] Douglas A. Gwyn. C standards update. The Journal of C Language Translation, 6(2):86–89, December 1994. ISSN 1042-5721.

Gwyn:JCLT-6-3-150

[Gwy95] Douglas A. Gwyn. C standards update. The Journal of C Language Translation, 6(3):150–156, March 1995. ISSN 1042-5721.

Hoffmann:JCLT-5-2-70

[Hof93] Ulrich Hoffmann. Using C as target code for high-level programming languages. The Journal of C Language Translation, 5(2):70–90, December 1993. ISSN 1042-5721.

Holly:JCLT-1-2-104

[Hol89] Mike Holly. A parallel processing implementation. The Journal of C Language Translation, 1(2):104– 111, September 1989. ISSN 1042-5721.

Homer:JCLT-4-3-209

[Hom93] Bill Homer. Aliasing control via restricted pointers. The Journal of C Language Translation, 4(3):209– 222, March 1993. ISSN 1042-5721.

Jaeschke:JCLT-0-1-32

[Jae89a] Rex Jaeschke. Author guidelines.

The Journal of C Language Translation, 0(1):32–38, March 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-2-96

[Jae89b] Rex Jaeschke. Electronic survey number 1. The Journal of C Language Translation, 1(2):96– 103, September 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-3-196

[Jae89c] Rex Jaeschke. Electronic survey number 2. The Journal of C Language Translation, 1(3):196–204, December 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-1-61

[Jae89d] Rex Jaeschke. Miscellanea. The Journal of C Language Translation, 1(1):61–67, June 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-2-147

[Jae89e] Rex Jaeschke. Miscellanea. The Journal of C Language Translation, 1(2):147–152, September 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-3-234

[Jae89f] Rex Jaeschke. Miscellanea. The Journal of C Language Translation, 1(3):234–244, December 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-2-82

[Jae89g] Rex Jaeschke. Numerical C extensions group. The Journal of C Language Translation, 1(2):82, September 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-3-180

[Jae89h] Rex Jaeschke. Numerical C extensions group. The Journal of C Language Translation, 1(3):180, December 1989. ISSN 1042-5721.

Jaeschke:JCLT-0-1-31

[Jae89i] Rex Jaeschke. Numerical C extensions group formed. The Journal of C Language Translation, 0(1): 31, March 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-1-9

[Jae89j] Rex Jaeschke. Numerical C Extensions Group status. The Journal of C Language Translation, 1(1): 9–11, June 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-2-115

[Jae89k] Rex Jaeschke. Parallel programming: Linda meets C, Part II - J. Leichter. The Journal of C Language Translation, 1(2):115–121, September 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-1-32

[Jae89l] Rex Jaeschke. Pragmania. The Journal of C Language Translation, 1(1):32–47, June 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-2-138

[Jae89m] Rex Jaeschke. Pragmania. The Journal of C Language Translation, 1(2):138–140, September 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-3-161

[Jae89n] Rex Jaeschke. Pragmania. The Journal of C Language Translation, 1(3):161–165, December 1989. ISSN 1042-5721.

Jaeschke:JCLT-0-1-17

[Jae89o] Rex Jaeschke. A prototype extraction tool. The Journal of C Language Translation, 0(1):17–30, March 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-2-127

[Jae89p] Rex Jaeschke. Validation suite report. The Journal of C Language Translation, 1(2):127–137, September 1989. ISSN 1042-5721.

Jaeschke:JCLT-1-4-275

[Jae90a] Rex Jaeschke. Electronic survey number 3. The Journal of C Language Translation, 1(4):275–282, March 1990. ISSN 1042-5721.

Jaeschke:JCLT-2-1-22

[Jae90b] Rex Jaeschke. Electronic survey number 3. The Journal of C Language Translation, 2(1):22–28, June 1990. ISSN 1042-5721.

Jaeschke:JCLT-2-2-95

[Jae90c] Rex Jaeschke. Electronic survey number 5. The Journal of C Language Translation, 2(2):95– 101, September 1990. ISSN 1042-5721.

Jaeschke:JCLT-2-3-185

[Jae90d] Rex Jaeschke. Electronic survey number 6. The Journal of C Language Translation, 2(3):185–193, December 1990. ISSN 1042-5721.

Jaeschke:JCLT-1-4-316

[Jae90e] Rex Jaeschke. Miscellanea. The Journal of C Language Translation, 1(4):316–332, March 1990. ISSN 1042-5721.

Jaeschke:JCLT-2-1-71

[Jae90f] Rex Jaeschke. Miscellanea. The Journal of C Language Translation, 2(1):71–80, June 1990. ISSN 1042-5721.

Jaeschke:JCLT-2-2-152

[Jae90g] Rex Jaeschke. Miscellanea. The Journal of C Language Translation, 2(2):152–164, September 1990. ISSN 1042-5721.

Jaeschke:JCLT-2-3-238

[Jae90h] Rex Jaeschke. Miscellanea. The Journal of C Language Translation, 2(3):238–248, December 1990. ISSN 1042-5721.

Jaeschke:JCLT-1-4-253

[Jae90i] Rex Jaeschke. Numerical C extensions group. The Journal of C Language Translation, 1(4):253– 254, March 1990. ISSN 1042-5721.

Jaeschke:JCLT-1-4-305

[Jae90j] Rex Jaeschke. Pragmania. The Journal of C Language Translation, 1(4):305–309, March 1990. ISSN 1042-5721.

Jaeschke:JCLT-2-1-57

[Jae90k] Rex Jaeschke. Pragmania. The Journal of C Language Translation, 2(1):57–60, June 1990. ISSN 1042-5721.

Jaeschke:JCLT-2-3-216

[Jae90l] Rex Jaeschke. Pragmania. The Journal of C Language Translation, 2(3):216–224, December 1990. ISSN 1042-5721.

Jaeschke:JCLT-3-3-198

[Jae91a] Rex Jaeschke. Electronic survey number 10. The Journal of C Language Translation, 3(3):198–205, December 1991. ISSN 1042-5721.

Jaeschke:JCLT-2-4-275

[Jae91b] Rex Jaeschke. Electronic survey number 7. The Journal of C Language Translation, 2(4):275–286, March 1991. ISSN 1042-5721.

Jaeschke:JCLT-3-1-62

[Jae91c] Rex Jaeschke. Electronic survey number 8. The Journal of C Language Translation, 3(1):62–69, June 1991. ISSN 1042-5721.

Jaeschke:JCLT-3-2-121

[Jae91d] Rex Jaeschke. Electronic survey number 9. The Journal of C Language Translation, 3(2):121–129, September 1991. ISSN 1042-5721.

Jaeschke:JCLT-2-4-332

[Jae91e] Rex Jaeschke. Miscellanea. The Journal of C Language Translation, 2(4):332–340, March 1991. ISSN 1042-5721.

Jaeschke:JCLT-3-1-81

[Jae91f] Rex Jaeschke. Miscellanea. The Journal of C Language Translation, 3(1):81–87, June 1991. ISSN 1042-5721.

Jaeschke:JCLT-3-2-162

[Jae91g] Rex Jaeschke. Miscellanea. The Journal of C Language Translation, 3(2):162–166, September 1991. ISSN 1042-5721.

Jaeschke:JCLT-3-3-248

[Jae91h] Rex Jaeschke. Miscellanea. The Journal of C Language Translation, 3(3):248–251, December 1991. ISSN 1042-5721.

Jaeschke:JCLT-3-2-154

[Jae91i] Rex Jaeschke. Pragmania. The Journal of C Language Translation, 3(2):154–155, September 1991. ISSN 1042-5721.

Jaeschke:JCLT-3-4-275

[Jae92a] Rex Jaeschke. Electronic survey number 11. The Journal of C Language Translation, 3(4):275–280, March 1992. ISSN 1042-5721.

Jaeschke:JCLT-3-4-338

[Jae92b] Rex Jaeschke. Miscellanea. The Journal of C Language Translation, 3(4):338–340, March 1992. ISSN 1042-5721.

Jaeschke:JCLT-3-4-291

[Jae92c] Rex Jaeschke. Pragmania. The Journal of C Language Translation, 3(4):291–294, March 1992. ISSN 1042-5721.

Jaffe:JCLT-2-1-61

[Jaf90] Martha Jaffe. A language compatible arithmetic standard. The Journal of C Language Translation, 2(1):61–70, June 1990. ISSN 1042-5721.

Jervis:JCLT-3-3-177

[Jer91] Bob Jervis. Parasol: A C-based systems programming language, Part I. The Journal of C Language Translation, 3(3):177–197, December 1991. ISSN 1042-5721.

Jervis:JCLT-3-4-319

[Jer92a] Bob Jervis. Parasol: A C-based systems programming language, Part II. The Journal of C Language Translation, 3(4):319–337, March 1992. ISSN 1042-5721.

Jervis:JCLT-4-1-27

[Jer92b] Robert Jervis. NCEG progress report. The Journal of C Language Translation, 4(1):27–41, September 1992. ISSN 1042-5721.

Jervis:JCLT-4-2-87

[Jer92c] Robert Jervis. An NCEG progress report: December, 1992. The Journal of C Language Translation, 4(2):87–89, December 1992. ISSN 1042-5721.

Kohlmiller:JCLT-3-1-38

[Koh91a] Paul Kohlmiller. C/C++ compatibility. The Journal of C Language Translation, 3(1):38–41, June 1991. ISSN 1042-5721.

Kohlmiller:JCLT-3-2-130

[Koh91b] Paul Kohlmiller. C/C++ compatibility. The Journal of C Language Translation, 3(2):130–139, September 1991. ISSN 1042-5721.

Kohlmiller:JCLT-3-3-244

[Koh91c] Paul Kohlmiller. C/C++ compatibility. The Journal of C Language Translation, 3(3):244–247, December 1991. ISSN 1042-5721.

Kurver:JCLT-2-4-267

[Kur91a] Rob E. H. Kurver. A parallel extension to ANSI C. The Journal of C Language Translation, 2(4):267– 274, March 1991. ISSN 1042-5721.

Kurver:JCLT-3-1-12

[Kur91b] Rob E. H. Kurver. A Standard C compiler for the transputer. The Journal of C Language Translation, 3(1):12–19, June 1991. ISSN 1042-5721.

Lally:JCLT-2-2-119

[Lal90] Steven Lally. Concurrent C: A language for multi-processing. The Journal of C Language Translation, 2(2):119–132, September 1990. ISSN 1042-5721.

Leary:JCLT-2-1-45

[Lea90] Kevin Leary. C and digital signal processing. The Journal of C Language Translation, 2(1):45–47, June 1990. ISSN 1042-5721.

Leichter:JCLT-1-3-181

[Lei89a] Leichter. Parallel programming:

Linda meets C, Part III. The Journal of C Language Translation, 1 (3):181–195, December 1989. ISSN 1042-5721.

Leichter:JCLT-1-1-55

[Lei89b] Jerry Leichter. Parallel programming: Linda meets C, Part I. The Journal of C Language Translation, 1(1):55–60, June 1989. ISSN 1042-5721.

Leichter:JCLT-1-4-283

[Lei90] Leichter. Parallel programming:

Linda meets C, Part IV. The Journal of C Language Translation, 1 (4):283–291, March 1990. ISSN 1042-5721.

Leichter:JCLT-2-4-296

[Lei91] Jerrold Leichter. Initializers and finalizers: A proposed C extension. The Journal of C Language Translation, 2(4):296–304, March 1991. ISSN 1042-5721.

Lenkov:JCLT-1-2-112

[Len89] Dmitry Lenkov. C++ standardization project report. The Journal of C Language Translation, 1(2):112–114, September 1989. ISSN 1042-5721.

Levine:JCLT-4-1-70

[Lev92a] John Levine. Miscellanea. The Journal of C Language Translation, 4(1):70–72, September 1992. ISSN 1042-5721.

Levine:JCLT-4-2-161

[Lev92b] John Levine. Miscellanea. The Journal of C Language Translation, 4(2):161–162, December 1992. ISSN 1042-5721.

Levine:JCLT-4-4-296

[Lev93a] John Levine. Electronic survey number 12. The Journal of C Language Translation, 4(4):296–300, June 1993. ISSN 1042-5721.

Levine:JCLT-5-2-103

[Lev93b] John Levine. Electronic survey number 13. The Journal of C Language Translation, 5(2):103–108, December 1993. ISSN 1042-5721.

Levine:JCLT-4-3-250

[Lev93c] John Levine. Miscellanea. The Journal of C Language Translation, 4(3):250, March 1993. ISSN 1042-5721.

Levine:JCLT-4-4-301

[Lev93d] John Levine. Miscellanea. The Journal of C Language Translation, 4(4):301–??, June 1993. ISSN 1042-5721.

Levine:JCLT-5-1-57

[Lev93e] John Levine. Miscellanea. The Journal of C Language Translation, 5(1):57–59, September 1993. ISSN 1042-5721.

Levine:JCLT-5-2-116

[Lev93f] John Levine. Miscellanea. The Journal of C Language Translation, 5(2):116–126, December 1993. ISSN 1042-5721.

Levine:JCLT-4-3-182

[Lev93g] John R. Levine. An NCEG progress report. The Journal of C Language Translation, 4(3):182– 184, March 1993. ISSN 1042-5721.

Levine:JCLT-5-2-109

[Lev93h] John R. Levine. Pragmania. The Journal of C Language Translation, 5(2):109–115, December 1993. ISSN 1042-5721.

Levine:JCLT-5-3-199

[Lev94a] John Levine. Miscellanea. The Journal of C Language Translation, 5(3):199–202, March 1994. ISSN 1042-5721.

Levine:JCLT-5-4-254

[Lev94b] John Levine. Miscellanea. The Journal of C Language Translation, 5(4):254–255, June 1994. ISSN 1042-5721.

Levine:JCLT-6-1-61

[Lev94c] John Levine. Miscellanea. The Journal of C Language Translation, 6(1):61–64, September 1994. ISSN 1042-5721.

Levine:JCLT-6-2-137

[Lev94d] John Levine. Miscellanea. The Journal of C Language Translation, 6(2):137–138, December 1994. ISSN 1042-5721.

Levine:JCLT-6-3-183

[Lev95] John Levine. Miscellanea. The Journal of C Language Translation, 6(3):183–186, March 1995. ISSN 1042-5721.

Long:JCLT-3-1-70

[Lon91] Paul Long. Emitting C source. The Journal of C Language Translation, 3(1):70–80, June 1991. ISSN 1042-5721.

Long:JCLT-6-3-157

[Lon95] Paul Long. A graphical model of the Standard C type system. The Journal of C Language Translation, 6(3):157–161, March 1995. ISSN 1042-5721.

Aki:JCLT-3-2-101

[MA91] W. M. McKeeman and Shota Aki. Reusable incremental scanning. The Journal of C Language Translation, 3(2):101–120, September 1991. ISSN 1042-5721.

McKeeman:JCLT-4-1-53

[MA92a] W. M. McKeeman and Shota Aki. Reusable incremental C preprocessing, Part I. The Journal of C Language Translation, 4(1):53–69, September 1992. ISSN 1042-5721.

McKeeman:JCLT-4-2-117

[MA92b] W. M. McKeeman and Shota Aki. Reusable incremental C preprocessing, Part II. The Journal of C Language Translation, 4(2):117– 125, December 1992. ISSN 1042-5721.

McKeeman:JCLT-4-3-191

[MA93] W. M. McKeeman and Shota Aki. Reusable incremental C parsing. The Journal of C Language Translation, 4(3):191–202, March 1993. ISSN 1042-5721.

McKeeman:JCLT-2-3-177

[MAA90] W. M. McKeeman, Shota Aki, and Scot Aurenz. Parser-independent compilers. The Journal of C Language Translation, 2(3):177–184, December 1990. ISSN 1042-5721.

MacDonald:JCLT-1-1-20

[Mac89a] Tom MacDonald. Adding complex arithmetic to C. The Journal of C Language Translation, 1(1):20–31, June 1989. ISSN 1042-5721.

MacDonald:JCLT-1-2-83

[Mac89b] Tom MacDonald. Aliasing issues in C. The Journal of C Language Translation, 1(2):83–95, September 1989. ISSN 1042-5721.

MacDonald:JCLT-0-1-9

[Mac89c] Tom MacDonald. C language and numerical programming. The Journal of C Language Translation, 0(1):9–16, March 1989. ISSN 1042-5721.

MacDonald:JCLT-1-3-215

[Mac89d] Tom MacDonald. Variable length arrays. The Journal of C Language Translation, 1(3):215–233, December 1989. ISSN 1042-5721.

MacDonald:JCLT-1-4-292

[Mac90a] Tom MacDonald. Complex and VLAs revisited. The Journal of C Language Translation, 1(4):292– 299, March 1990. ISSN 1042-5721.

MacDonald:JCLT-1-4-310

[Mac90b] Tom MacDonald. Cray C: C in a 64-bit world. The Journal of C Language Translation, 1(4):310– 315, March 1990. ISSN 1042-5721.

MacDonald:JCLT-2-2-102

[Mac90c] Tom MacDonald. IEEE floatingpoint arithmetic and C. The Journal of C Language Translation, 2(2):102–112, September 1990. ISSN 1042-5721.

MacDonald:JCLT-2-1-9

[Mac90d] Tom MacDonald. Parsing initializers. The Journal of C Language Translation, 2(1):9–21, June 1990. ISSN 1042-5721.

MacDonald:JCLT-2-3-225

[Mac90e] Tom MacDonald. Restricted pointers. The Journal of C Language Translation, 2(3):225–237, December 1990. ISSN 1042-5721.

MacDonald:JCLT-3-2-140

[Mac91a] Tom MacDonald. Adding a bit type to C. The Journal of C Language Translation, 3(2):140–153, September 1991. ISSN 1042-5721.

MacDonald:JCLT-2-4-305

[Mac91b] Tom MacDonald. Cray C and Fortran interlanguage communication. The Journal of C Language Translation, 2(4):305–317, March 1991. ISSN 1042-5721.

MacDonald:JCLT-3-3-229

[Mac91c] Tom MacDonald. Massively parallel C: Architectures and data distribution. The Journal of C Language Translation, 3(3):229–243, December 1991. ISSN 1042-5721.

MacDonald:JCLT-3-1-20

[Mac91d] Tom MacDonald. NCEG progress report. The Journal of C Language Translation, 3(1):20–37, June 1991. ISSN 1042-5721.

MacDonald:JCLT-3-4-281

[Mac92] Tom MacDonald. Massively parallel C: Data parallel execution. The Journal of C Language Translation, 3(4):281–290, March 1992. ISSN 1042-5721.

Martin:JCLT-1-2-141

[Mar89] N. Martin. European conformance testing service for C. The Journal of C Language Translation, 1(2): 141–146, September 1989. ISSN 1042-5721.

Martin:JCLT-3-3-227

[Mar91] Neil Martin. European C conformance testing. The Journal of C Language Translation, 3(3):227– 228, December 1991. ISSN 1042-5721.

Martin:JCLT-2-3-173

[MC90] Neil Martin and Dan Chacon.

European C conformance testing. The Journal of C Language Translation, 2(3):173–176, December 1990. ISSN 1042-5721.

McKeeman:JCLT-3-3-214

[McK91a] W. M. McKeeman. Cb (C flat): A low-level subset of C. The Journal of C Language Translation, 3 (3):214–226, December 1991. ISSN 1042-5721.

McKeeman:JCLT-2-4-259

[McK91b] W. M. McKeeman. Resolving typedefs in a multipass C compiler. The Journal of C Language Translation, 2(4):259–266, March 1991. ISSN 1042-5721.

McKeeman:JCLT-4-4-291

[McK93a] W. M. McKeeman. Avoiding arithmetic exceptions in constant expression evaluation. The Journal of C Language Translation, 4(4):291–295, June 1993. ISSN 1042-5721.

McKeeman:JCLT-5-1-27

[McK93b] W. M. McKeeman. Avoiding overflow in constant expression evaluation. The Journal of C Language Translation, 5(1):27–31, September 1993. ISSN 1042-5721.

Meloy:JCLT-1-1-12

[Mel89] Sue Meloy. A solution to name space pollution. The Journal of C Language Translation, 1(1):12–16, June 1989. ISSN 1042-5721.

Micco:JCLT-5-3-139

[Mic94a] John Micco. Standardizing C++: 1993 part 1. The Journal of C Language Translation, 5(3):139–162, March 1994. ISSN 1042-5721.

Micco:JCLT-5-4-212

[Mic94b] John Micco. Standardizing C++: 1993 part 2. The Journal of C Language Translation, 5(4):212–228, June 1994. ISSN 1042-5721.

Micco:JCLT-6-1-10

[Mic94c] John Micco. Standardizing C++: 1994 part 1. The Journal of C Language Translation, 6(1):10–19, September 1994. ISSN 1042-5721.

Micco:JCLT-6-2-73

[Mic94d] John Micco. Standardizing C++: 1994 part 2. The Journal of C Language Translation, 6(2):73–85, December 1994. ISSN 1042-5721.

Micco:JCLT-6-3-162

[Mic95] John Micco. Standardizing C++:

1994 part 3. The Journal of C Language Translation, 6(3):162–178, March 1995. ISSN 1042-5721.

Parks:JCLT-2-3-194

[Par90] John H. Parks. Case study: Building an ANSI CPP. The Journal of C Language Translation, 2(3):194– 206, December 1990. ISSN 1042-5721.

Pennello:JCLT-2-4-318

[Pen91] Thomas J. Pennello. Iterators.

The Journal of C Language Translation, 2(4):318–331, March 1991. ISSN 1042-5721.

Plauger:JCLT-1-3-153

[Pla89a] P. J. Plauger. Diagnosing quiet changes. The Journal of C Language Translation, 1(3):153–160, December 1989. ISSN 1042-5721.

Plauger:JCLT-0-1-1

[Pla89b] P. J. Plauger. Internationalizing C. The Journal of C Language Translation, 0(1):1–8, March 1989. ISSN 1042-5721.

Plauger:JCLT-1-1-1

[Pla89c] P. J. Plauger. Pointers and addresses. The Journal of C Language Translation, 1(1):1–8, June 1989. ISSN 1042-5721.

Plauger:JCLT-1-2-73

[Pla89d] P. J. Plauger. Translating Pascal to C. The Journal of C Language Translation, 1(2):73–81, September 1989. ISSN 1042-5721.

Plauger:JCLT-2-3-165

[Pla90a] P. J. Plauger. Generalizing type qualifiers. The Journal of C Language Translation, 2(3):165–172, December 1990. ISSN 1042-5721.

Plauger:JCLT-2-1-1

[Pla90b] P. J. Plauger. Sequence points and parallelism. The Journal of C Language Translation, 2(1):1–8, June 1990. ISSN 1042-5721.

Plauger:JCLT-1-4-245

[Pla90c] P. J. Plauger. Standard C library name space control. The Journal of C Language Translation, 1(4): 245–252, March 1990. ISSN 1042-5721.

Plauger:JCLT-2-2-87

[Pla90d] P. J. Plauger. Standard C: What barely works. The Journal of C Language Translation, 2(2):87–94, September 1990. ISSN 1042-5721.

Plauger:JCLT-3-3-169

[Pla91a] P. J. Plauger. Character sets and C. The Journal of C Language Translation, 3(3):169–176, December 1991. ISSN 1042-5721.

Plauger:JCLT-3-2-89

[Pla91b] P. J. Plauger. Floating-point primitives. The Journal of C Language Translation, 3(2):89–100, September 1991. ISSN 1042-5721.

Plauger:JCLT-2-4-249

[Pla91c] P. J. Plauger. Implementing locales. The Journal of C Language Translation, 2(4):249–258, March 1991. ISSN 1042-5721.

Plauger:JCLT-3-1-1

[Pla91d] P. J. Plauger. Translating multibyte characters. The Journal of C Language Translation, 3(1):1– 11, June 1991. ISSN 1042-5721.

Plauger:JCLT-3-4-253

[Pla92a] P. J. Plauger. A C++ library.

The Journal of C Language Translation, 3(4):253–258, March 1992. ISSN 1042-5721.

Plauger:JCLT-4-1-1

[Pla92b] P. J. Plauger. Internationalization.

The Journal of C Language Translation, 4(1):1–8, September 1992. ISSN 1042-5721.

Plauger:JCLT-4-2-73

[Pla92c] P. J. Plauger. Turtles and overcoats. The Journal of C Language Translation, 4(2):73–79, December 1992. ISSN 1042-5721.

Plauger:JCLT-4-4-251

[Pla93a] P. J. Plauger. Interpreting the C Standard. The Journal of C Language Translation, 4(4):251–263, June 1993. ISSN 1042-5721.

Plauger:JCLT-4-3-163

[Pla93b] P. J. Plauger. The normative addendum to Standard C. The Journal of C Language Translation, 4 (3):163–170, March 1993. ISSN 1042-5721.

Plauger:JCLT-5-2-61

[Pla93c] P. J. Plauger. The Standard C++ library. The Journal of C Language Translation, 5(2):61–69, December 1993. ISSN 1042-5721.

Plauger:JCLT-5-1-1

[Pla93d] P. J. Plauger. Wide character iostreams. The Journal of C Language Translation, 5(1):1–10, September 1993. ISSN 1042-5721.

Plauger:JCLT-6-1-1

[Pla94a] P. J. Plauger. Managing the heap. The Journal of C Language Translation, 6(1):1–9, September 1994. ISSN 1042-5721.

Plauger:JCLT-5-4-203

[Pla94b] P. J. Plauger. Math in all precisions. The Journal of C Language Translation, 5(4):203–211, June 1994. ISSN 1042-5721.

Plauger:JCLT-5-3-127

[Pla94c] P. J. Plauger. Revising Standard C. The Journal of C Language Translation, 5(3):127–133, March 1994. ISSN 1042-5721.

Plauger:JCLT-6-2-65

[Pla94d] P. J. Plauger. What’s in a byte? The Journal of C Language Translation, 6(2):65–72, December 1994. ISSN 1042-5721.

Plauger:JCLT-6-3-139

[Pla95] P. J. Plauger. Templates in C++. The Journal of C Language Translation, 6(3):139–149, March 1995. ISSN 1042-5721.

Pavlov:JCLT-4-2-142

[PM92] V. S. Pavlov and S. A. Mironov. A universal C compiler. The Journal of C Language Translation, 4 (2):142–160, December 1992. ISSN 1042-5721.

Prosser:JCLT-4-3-203

[Pro93] David F. Prosser. Designated initializers and compound literals. The Journal of C Language Translation, 4(3):203–208, March 1993. ISSN 1042-5721.

Ritchie:JCLT-2-2-81

[Rit90] Dennis Ritchie. Variable-size arrays in C. The Journal of C Language Translation, 2(2):81–86, September 1990. ISSN 1042-5721.

Saks:JCLT-4-1-17

[Sak92] Dan Saks. Standardizing C++:

The first meetings. The Journal of C Language Translation, 4(1):17– 26, September 1992. ISSN 1042-5721.

Saks:JCLT-4-3-171

[Sak93a] Dan Saks. Standardizing C++:

The first year of joint meetings. The Journal of C Language Translation, 4(3):171–181, March 1993. ISSN 1042-5721.

Saks:JCLT-4-4-264

[Sak93b] Dan Saks. Standardizing C++:

The rest of 1992, part 1. The Journal of C Language Translation, 4(4):264–276, June 1993. ISSN 1042-5721.

Saks:JCLT-5-1-11

[Sak93c] Dan Saks. Standardizing C++:

The rest of 1992, part 2. The Journal of C Language Translation, 5 (1):11–17, September 1993. ISSN 1042-5721.

Spinellis:JCLT-4-3-238

[Spi93] Diomidis D. Spinellis. Checking C declarations at link time. The Journal of C Language Translation, 4(3):238–249, March 1993. ISSN 1042-5721.

Stanberry:JCLT-4-4-284

[Sta93a] Linda Stanberry. NCEG progress report. The Journal of C Language Translation, 4(4):284–290, June 1993. ISSN 1042-5721.

Stanberry:JCLT-5-1-22

[Sta93b] Linda Stanberry. NCEG progress report. The Journal of C Language Translation, 5(1):22–26, September 1993. ISSN 1042-5721.

Stanberry:JCLT-5-2-95

[Sta93c] Linda Stanberry. NCEG progress report. The Journal of C Language Translation, 5(2):95–102, December 1993. ISSN 1042-5721.

Stanberry:JCLT-6-1-35

[Sta94a] Linda Stanberry. DPCE status report. The Journal of C Language Translation, 6(1):35–43, September 1994. ISSN 1042-5721.

Stanberry:JCLT-5-3-175

[Sta94b] Linda Stanberry. NCEG progress report. The Journal of C Language Translation, 5(3):175–182, March 1994. ISSN 1042-5721.

Stanberry:JCLT-5-4-246

[Sta94c] Linda Stanberry. NCEG progress report. The Journal of C Language Translation, 5(4):246–253, June 1994. ISSN 1042-5721.

Stanberry:JCLT-6-2-90

[Sta94d] Linda Stanberry. NCEG progress report. The Journal of C Language Translation, 6(2):90–99, December 1994. ISSN 1042-5721.

Sterling:JCLT-2-2-113

[Ste90] Gordon A. Sterling. Signal handling and C. The Journal of C Language Translation, 2(2):113– 118, September 1990. ISSN 1042-5721.

Coonen:JCLT-5-3-134

[TC94] Jim Thomas and Jerome T. Coonen. Issues regarding imaginary types for C and C++. The Journal of C Language Translation, 5(3): 134–138, March 1994. ISSN 1042-5721.

Tydeman:JCLT-4-2-90

[Tyd92] Fred Tydeman. Extending C for numerical computing (NCEG). The Journal of C Language Translation, 4(2):90–116, December 1992. ISSN 1042-5721.

Tydeman:JCLT-6-3-179

[Tyd95] Fred Tydeman. Numerical issues.

The Journal of C Language Translation, 6(3):179–182, March 1995. ISSN 1042-5721.